OpenStudy (anonymous):

x-2y=-8? (finding slope)

OpenStudy (anonymous):

need some help?

OpenStudy (anonymous):

alright the easiest way to find the slope is to convert it to slope intercept form

OpenStudy (anonymous):

sorry i gotta go,,, remember y=mx+b, m is the slope

OpenStudy (anonymous):

every time I "finish" my equation it becomes wrong somehow. so far I've done -2Y=-8-x and divide that which gets me Y=4-x

OpenStudy (anonymous):

they're saying it's wrong...

OpenStudy (firekat97):

x - 2y = -8 so you want to get it in the form y = mx + b so you can simply read off the slope

OpenStudy (firekat97):

so try rearranging the equation and post up what you get

OpenStudy (anonymous):

slope equals positive 1/2

OpenStudy (whpalmer4):

\[x-2y=-8\] One way to get the slope (as other posters have suggested) is to rearrange the equation in slope-intercept form, or \[y = mx + b\]You can then "read off" the slope (\(m\)) and the y-intercept value (\(b\)). To do this, just solve your equation for \(y\): \[x-2y = -8\]add \(2y\) to both sides \[x-2y+2y = -8+2y\]\[x=-8+2y\]add \(8\) to both sides \[x+8 = -8+8+2y\]\[x+8 = 2y\]divide both sides by \(2\)\[\frac{x}{2} + \frac{8}{2} = \frac{2y}{2}\]\[\frac{1}{2} x + 4 = y\]swap sides for conveniences \[y = \frac{1}{2} x + 4\] comparing with our slope-intercept form \(y = mx + b\) we see that \(m =\frac{1}{2}\) and so our slope is \(\frac{1}{2}\).

OpenStudy (whpalmer4):

Another approach is to just find two points on the line and use the formula for the slope of a line passing through two points: \[m = \frac{y_2-y_1}{x_2-x_1}\]where the line passes through points \((x_1,y_1)\) and \((x_2,y_2)\) To find out points, pick some values that will be easy to compute. \(x=0\) is usually a good one, let's try it: \[x=0,\ y = \frac{0}{2}x+4= 4\]so the point \((0,2)\) is our first point, \((x_1,y_1)\) As we have that pesky \(\frac{1}{2}\) in the formula, let's use a value of \(2\) for \(x\) for the other point to eliminate the fraction: \[x = 2, \ y = \frac{1}{2}(2) + 4 = 5\] So \((2,5)\) is our other point, \((x_2,y_2)\) Let's plug those points into the formula for the slope: \[m = \frac{y_2-y_1}{x_2-x_1} = \frac{5-4}{2-0} = \frac{1}{2}\] Look at that, we got the same answer. Good thing, too!

OpenStudy (whpalmer4):

sorry, I mistyped that equation for the first point in the previous post: \[x=0,\ y = \frac{0}{2}x+4= 4\]should be\[x=0,\ y = \frac{1}{2}(0)+4= 4\]